Exercise 1.47 Find the angle between each of tha following pairs of vectors Ari Ag, j and B Bri By j Part A Azu 2.00, A. 5.20 B 1.10, B -3.80. Submit My Answers Give Up Part B A2 3.20, An 4.20, Ba 11-8, BV, 6.20 Submit My Answers Give Up Part C -4.00, Ath, 2.00, B 7.00 B 4.00 Submit My Answers Give Up
Answer #1

As we know that- Cos\Theta = \underset{(A}{\rightarrow}\underset{.B)}{\rightarrow} / (\left | A \right |.\left | B \right |)

For Part a)- \underset{(A}{\rightarrow}\underset{.B)}{\rightarrow} = (-2 )(1.10 ) + ( 5.20)(-3.80 ) = -21.96

And (\left | A \right |.\left | B \right |) = (5.57)( 3.96) = 22.06

so Cos\Theta = -21.96 / 22.06 = -0.89

than \Theta = cos-1( -0.89) = 152.87o

Now the same formula will use for the rest of the parts of question.

Part b)- Cos\Theta = 63.80 / 70.38 = 0.91

So \Theta = 24.97o

Part C)-  Cos\Theta = 0

so, \Theta = 90o

Thank you !!! Have fun